# The Leftmost particle of branching subordinators

### v1

@@ Appendix A Construction of branching subordinators
Our construction borrows most of its arguments from [22] and [13] but requires a slight adaptation due to the fact that we relax the classical exponential assumption (6), so we shall not give all details and we may refer to original arguments when needed. Before getting started, let us recall a few notations and definitions. The set
@@
stands for the set of sequences of $\mathscr{Q}$ with at least two finite terms and we assume that $\Lambda(\mathscr{Q}^{\star})\in(0,\infty]$, see (3). Recall (see equation (4)) that the integral $\int_{\mathscr{Q}}\big{|}\sharp\{n\geq 1:\;x_{n}<a\}-1\big{|}\Lambda(\mathrm{d}\bm{x})$ is finite for any $a>0$. This hypothesis aims to replace the classical exponential assumption (6) and is enough to ensure that for any $a>0$ and $\lambda\geq 0$
@@
Also recall that $T_{a}:\bm{x}\in\mathscr{Q}\mapsto\bm{x}^{a}\in\mathscr{Q}$ is defined by $(\bm{x}^{a})_{1}=x_{1}$ and for any $i\geq 2$, $(\bm{x}^{a})_{i}=x_{i}\mathds{1}_{\{x_{i}<a\}}+\infty\mathds{1}_{\{x_{i}\geq a\}}$, and that $\Lambda^{a}(\mathrm{d}\bm{x})=\mathds{1}_{\{\bm{x}\neq(0,\emptyset)\}}(T_{a}\Lambda)(\mathrm{d}\bm{x})$, where $T_{a}\Lambda$ denotes the image measure of $\Lambda$ by $T_{a}$. We see from the definition of $\Lambda^{a}$ and (4) that $\Lambda^{a}$ is supported on $\{\bm{x}\in\mathscr{Q};\;\sharp\{i\geq 2,x_{i}<\infty\}<\infty\}$. Recall that $\lim_{a\to\infty}\Lambda^{a}(\mathscr{Q}^{\star})>0$ (see equation (3)) so there exists $a_{0}(\Lambda)\geq 0$ such that $\Lambda^{a}(\mathscr{Q}^{\star})>0$ for all $a>a_{0}(\Lambda)$. On the other hand, $\Lambda^{a}(\mathscr{Q}^{\star})\leq M_{a}(0)<\infty$. In particular, the measure $\Lambda^{a,\star}:=\Lambda^{a}(\cdot\cap\mathscr{Q}^{\star})/\Lambda^{a}(\mathscr{Q}^{\star})$ defined on $\mathscr{Q}^{\star}$ is a probability measure for all $a>a_{0}(\Lambda)$. From now one we always consider such an $a$. Now, let us introduce
@@
We are now ready to construct our branching subordinators. Denote by $\mathscr{U}:=\bigcup_{n\geq 0}((\mathbb{N}^{*})^{3})^{n}$ the set of finite sequences of $\mathbb{N}^{*}$-valued $3$-tuples with the convention $((\mathbb{N}^{*})^{3})^{0}=\{\varnothing\}$, where $\varnothing$ is the empty sequence. For a given $u\in\mathscr{U}$, there exists $n\in\mathbb{N}$ such that $u\in((\mathbb{N}^{*})^{3})^{n}$ and $n$ stands for the generation of the vertex $u$, denoted by $|u|$. In particular, the generation 00 only contains $\varnothing$. Any child of $u$ is denoted by $u\bm{p}$ instead of $(u,\bm{p})$ for some $\bm{p}\in(\mathbb{N}^{*})^{3}$. The tree $\mathscr{U}$ will play the role of indexing set for the particles of the branching subordinator $Y^{a}$ that we are about to construct. Precisely, for the child $u\bm{p}$ of $u$ with $\bm{p}=(q,k,i)$, the number $q$ will denote the level of this child, that is $q=1+\lfloor d_{u\bm{p}}\rfloor$ where $d_{u\bm{p}}$ stands for the distance between the birth position of the particle $u\bm{p}$ and the pre-branching position of the particle $u$. The number $k$ will mean that the particle $u\bm{p}$ was born during the $k$-th branching of the particle $u$ such that there exists a child $v$ of particle $u$ with level $q$ and the number $i$ will mean that particle $u\bm{p}$ is the $i$-th child of particle $u$ (that is the child with $i$-th smallest birth position) with level $q$ born during the latter branching. We note that some indices $u\in\mathscr{U}$ will possibly correspond to particles that are never born. An important aspect of the construction of $Y^{a}$ is that, almost surely, for all $u\in\mathscr{U}$, the sequence of branching times of the particle $u$ will be discrete, thus making possible the indexing of particles by the tree $\mathscr{U}$. That property is not true in general for the branching subordinator $Y$, which will therefore be constructed from $Y^{a}$ by a limiting procedure. We now proceed with the construction.
@@
Let $a>a_{0}(\Lambda)$ be a positive integer and introduce three independent families of random variables:
@@
  • $(T_{u,j}-T_{u,j-1})_{u\in\mathscr{U},j\geq 1}$ a family of i.i.d exponential random variables with mean $1/\Lambda^{a}(\mathscr{Q}^{\star})$ and $T_{u,0}=0$ for all $u\in\mathscr{U}$;
  • $(\Delta_{u,j})_{u\in\mathscr{U},j\geq 1}$ a family of i.i.d $\mathscr{Q}^{\star}$-valued random variables with law $\Lambda^{a,\star}$. The $i$-th term of the sequence $\Delta_{u,j}$ is denoted by $\Delta_{u,j}(i)$.
Note that for any $u\in\mathscr{U}$, $\sum_{j\geq 1}\delta_{T_{u,j}}$ is a Poisson point measure with intensity $\Lambda^{a}(\mathscr{Q}^{\star})\mathrm{d}t$. Let $N_{u}(t):=\sum_{j\geq 1}\mathds{1}_{\{T_{u,j}\leq t\}}$. We first deal with the initial particle $\varnothing$. The birth-time $\mathfrak{b}^{a}_{\varnothing}$ of particle $\varnothing$ equals 00 and for any $t\geq 0$,
$Y^{a}_{\varnothing}(t):=\xi_{\varnothing}(t)+\sum_{j=1}^{N_{\varnothing}(t)}\Delta_{\varnothing,j}(1),$ (44)
thus giving the generation 00. We now deal with the other particles. For that, let $q$ be a positive integer, $u\in\mathscr{U}$ with $u\not=\varnothing$ and define a new element $\Delta^{q}_{u,j}:=(\Delta^{q}_{u,j}(i))_{i\geq 1}\in\mathscr{Q}$ as follows: $\Delta^{q}_{u,j}=\emptyset$ if the set $\{l\geq 2;\;\Delta_{u,j}(l)\in[q-1,q)\}$ is empty. Otherwise, let $\tau^{q,-}_{u,j}:=\inf\{l\geq 2;\;\Delta_{u,j}(l)\in[q-1,q)\}$ and $\tau^{q,+}_{u,j}:=\inf\{l>2;\;\Delta_{u,j}(l)\geq q\}$, which is almost surely finite by definition. We set $\Delta^{q}_{u,j}(i):=\Delta_{u,j}(i+\tau^{q,-}_{u,j}-1)$ if $i\in\{1,\ldots,\tau^{q,+}_{u,j}-\tau^{q,-}_{u,j}\}$ and $\Delta^{q}_{u,j}(i):=\infty$ if $i>\tau^{q,+}_{u,j}-\tau^{q,-}_{u,j}$. One can see that $\Delta^{q}_{u,j}=\emptyset$ for all $q>a$.
@@
Now, let $n\in\mathbb{N}$, assume that $\{(u,Y^{a}_{u});\;|u|\leq n\}$ has been built and denote by $\mathfrak{b}^{a}_{u}$ the birth-time of any particle $u$ in generation $n$. Let $\bm{p}=(q,k,i)\in\mathbb{N}^{3}$. The birth-time $\mathfrak{b}^{a}_{u\bm{p}}$ of the child $u\bm{p}$ of particle $u$ is set to be $\infty$ if $\mathfrak{b}^{a}_{u}=\infty$, $B^{q}_{u}(k)=\infty$ or $\Delta^{q}_{u,B^{q}_{u}(k)}(i)=\infty$ and we set $Y^{a}_{u\bm{p}}=\infty$. In particular, $Y^{a}_{u\bm{p}}=\infty$ when $q>a$. Otherwise, it is given by
$\mathfrak{b}^{a}_{u\bm{p}}:=\mathfrak{b}^{a}_{u}+T_{u,B^{q}_{u}(k)}.$
@@
$Y^{a}_{u\bm{p}}(t):=Y^{a}_{u}\big{(}\mathfrak{b}^{a}_{u\bm{p}}-\big{)}+\Delta^{q}_{u,B^{q}_{u}(k)}(i)+\xi_{u\bm{p}}(t-\mathfrak{b}^{a}_{u\bm{p}})+\sum_{j=1}^{N_{u\bm{p}}(t-\mathfrak{b}^{a}_{u\bm{p}})}\Delta_{u\bm{p},j}(1).$ (45)
@@
Let $u=u_{1}\cdots u_{\mathfrak{j}}\in\mathscr{U}$ for some $\mathfrak{j}\geq 1$. We extend the definition of $Y^{a}_{u}$ on $[0,\mathfrak{b}^{a}_{u})$ in the following way: if $\mathfrak{j}\geq 2$, then for any $i\in\{1,\ldots,\mathfrak{j}-1\}$, $Y^{a}_{u}(t)=Y^{a}_{u_{1}\ldots u_{\mathfrak{i}}}(t)$ if $t\in[\mathfrak{b}^{a}_{u_{1}\cdots u_{\mathfrak{i}}},\mathfrak{b}^{a}_{u_{1}\cdots u_{\mathfrak{i}+1}})$ and $Y^{a}_{u}(t)=Y^{a}_{\varnothing}(t)$ if $t\in[0,\mathfrak{b}^{a}_{u_{1}})$. Otherwise, that is $\mathfrak{j}=1$, we set $Y^{a}_{u}(t)=Y^{q}_{\varnothing}(t)$ for any $t\in[0,\mathfrak{b}^{a}_{u})$. In other words, if the particle $u$ is not born yet at time $t$, then $Y^{a}_{u}(t)$ stands for the position at time $t$ of the most recent ancestor alive (at time $t$).
@@
$Y^{a}$ is referred to as a $a$-truncated branching subordinator. By a slight adaptation of Theorem 1.1(ii) in [13] we get that, for any $\lambda\geq 0$,
$\mathbb{E}\Big{[}\sum_{u\in\mathscr{N}^{a}(t)}e^{-\lambda Y^{a}_{u}(t)}\Big{]}=e^{-t\kappa_{a}(\lambda)},$ (46)
and $Y^{a}$ is a branching subordinator with finite birth intensity and characteristics $(d,\Lambda^{a})$ in the sense of Section 4.2 of [13].
@@
the canonical trajectory of particle $u$. Let $\bm{\xi}^{a}_{u}$, $u\in\mathscr{U}$ such that $\mathfrak{b}^{a}_{u}=\infty$, be i.i.d copies of $Y^{a}_{\varnothing}$ and independent of $(\bm{\xi}^{a}_{u};\;u\in\mathscr{U};\;\mathfrak{b}^{a}_{u}<\infty)$. Hence it is easily seen from (44) and (45) that, for all $a>a_{0}(\Lambda)$, $(\bm{\xi}^{a}_{u})_{u\in\mathscr{U}}$ is a collection of i.i.d subordinators with Laplace exponent $\phi$ and starting position 00.
@@
There exists a sequence $\{Y^{a},a\geq\lfloor a_{0}(\Lambda)\rfloor+1\}$ such that, for any $a\geq\lfloor a_{0}(\Lambda)\rfloor+1$, $Y^{a}$ is a branching subordinator with finite birth intensity $(d,\Lambda^{a})$ and, for any integer $a^{\prime}\geq a$
@@
We refer to Lemma 2.3 in [22] for a proof. For any $t\geq 0$, $\{Y^{a}(t),a\geq\lfloor a_{0}(\Lambda)\rfloor+1\}$ is an non-decreasing sequence of elements of $\mathscr{Q}$, which we recall is a subset of the space of point measures on $[0,\infty)$.
@@
We end this appendix with a technical lemma that will allow us to define without ambiguity the leftmost particle within some sets.
@@
Let $\Lambda$ be a measure on $\mathscr{Q}$ satisfying (3), (4) and Assumption 1 for some $\alpha\in(0,1)$. Let also $a>a_{0}(\Lambda)$. Then almost surely, in the branching subordinator $Y^{a}$ with characteristic couple $(0,\Lambda^{a})$, at all rational times $t$ (resp. at all times $t$ where a particle in $\mathscr{N}^{a}(t-)$ experiences a branching in $\mathscr{Q}^{\star}$, see (41)), there are finitely many particles in the system at time $t$ (resp. at time $t-$) and their positions are distinct.

### v2

@@ Appendix A Construction of branching subordinators
Our construction borrows most of its arguments from [26] and [14] but we slightly adapt their construction so that it fits our needs. We shall not give all details and we may refer to original arguments when needed. Before getting started, let us recall a few notations and definitions. The set
@@
stands for the set of sequences of $\mathscr{Q}$ with at least two finite terms and we assume that $\Lambda(\mathscr{Q}^{\star})\in(0,\infty]$, see (3). Recall (see equation (4)) that the integral $\int_{\mathscr{Q}}\big|\sharp\{n\geq 1;\;x_{n}<a\}-1\big|\Lambda(\mathrm{d}\bm{x})$ is finite for any $a>0$. This hypothesis aims to replace the classical exponential assumption (6) and is enough to ensure that for any $a>0$ and $\lambda\geq 0$
@@
Also recall that $T_{a}:\bm{x}\in\mathscr{Q}\mapsto\bm{x}^{a}\in\mathscr{Q}$ is defined by $(\bm{x}^{a})_{1}=x_{1}$ and for any $i\geq 2$, $(\bm{x}^{a})_{i}=x_{i}\mathds{1}_{\{x_{i}<a\}}+\infty\mathds{1}_{\{x_{i}\geq a\}}$, and that $\Lambda^{a}(\mathrm{d}\bm{x})=\mathds{1}_{\{\bm{x}\neq(0,\emptyset)\}}(T_{a}\Lambda)(\mathrm{d}\bm{x})$, where $T_{a}\Lambda$ denotes the image measure of $\Lambda$ by $T_{a}$. We see from the definition of $\Lambda^{a}$ and (4) that $\Lambda^{a}$ is supported on $\{\bm{x}\in\mathscr{Q};\;\sharp\{i\geq 2;x_{i}<\infty\}<\infty\}$. Recall that $\Lambda^{a}(\mathscr{Q}^{\star})=\Lambda(\{\bm{x}\in\mathscr{Q};x_{2}<a\})$ so $\lim_{a\to\infty}\Lambda^{a}(\mathscr{Q}^{\star})=\Lambda(\mathscr{Q}^{\star})>0$ (see equation (3)) so there exists $a_{0}(\Lambda)\geq 0$ such that $\Lambda^{a}(\mathscr{Q}^{\star})>0$ for all $a>a_{0}(\Lambda)$. On the other hand, $\Lambda^{a}(\mathscr{Q}^{\star})\leq M_{a}(0)<\infty$. In particular, the measure $\Lambda^{a,\star}:=\Lambda^{a}(\cdot\cap\mathscr{Q}^{\star})/\Lambda^{a}(\mathscr{Q}^{\star})$ defined on $\mathscr{Q}^{\star}$ is a probability measure for all $a>a_{0}(\Lambda)$. From now one we always consider such an $a$. Now, let us introduce
@@
We are now ready to construct our branching subordinators. Let $\mathbb{N}^{*}:=\{1,2,\ldots\}$ and denote by $\mathscr{U}:=\bigcup_{n\geq 0}((\mathbb{N}^{*})^{3})^{n}$ the set of finite sequences of $\mathbb{N}^{*}$-valued $3$-tuples with the convention $((\mathbb{N}^{*})^{3})^{0}=\{\varnothing\}$, where $\varnothing$ is the empty sequence. For a given $u\in\mathscr{U}$, there exists $n\geq 0$ such that $u\in((\mathbb{N}^{*})^{3})^{n}$ and $n$ stands for the generation of the vertex $u$, denoted by $|u|$. In particular, the generation 0 only contains $\varnothing$. Any child of $u$ is denoted by $u\bm{p}$ instead of $(u,\bm{p})$ for some $\bm{p}\in(\mathbb{N}^{*})^{3}$. The tree $\mathscr{U}$ will play the role of indexing set for the particles of the branching subordinator $Y^{a}$ that we are about to construct. Precisely, for the child $u\bm{p}$ of $u$ with $\bm{p}=(q,k,i)$, the number $q$ will denote the level of this child, that is $q=1+\lfloor d_{u\bm{p}}\rfloor$ where $d_{u\bm{p}}$ stands for the distance between the birth position of the particle $u\bm{p}$ and the pre-branching position of the particle $u$. The number $k$ will mean that the particle $u\bm{p}$ was born during the $k$-th branching of the particle $u$ such that there exists a child $v$ of particle $u$ with level $q$ and the number $i$ will mean that particle $u\bm{p}$ is the $i$-th child of particle $u$ (that is the child with $i$-th smallest birth position) with level $q$ born during the latter branching. We note that some indices $u\in\mathscr{U}$ will possibly correspond to particles that are never born. An important aspect of the construction of $Y^{a}$ is that, almost surely, for all $u\in\mathscr{U}$, the sequence of branching times of the particle $u$ will be discrete, thus making possible the indexing of particles by the tree $\mathscr{U}$. That property is not true in general for the branching subordinator $Y$, which will therefore be constructed from $Y^{a}$ by a limiting procedure. We now proceed with the construction.
@@
Let $a>a_{0}(\Lambda)$ be a positive integer and introduce two independent families of random variables:
@@
  • $(\mathscr{P}_{u})_{u\in\mathscr{U}}$, a family of i.i.d copies of a Poisson point process on $0,\infty)\times\mathscr{Q}^{\star}$ with intensity $\mathrm{d}t\otimes\Lambda^{a}(\cdot\cap\mathscr{Q}^{\star})$. For any borelian set $A\subset[0,\infty)$, introduce $N_{u}(A):=\sum_{\mathfrak{z}\in\mathscr{P}_{u}}\mathds{1}_{\{\mathfrak{z}\in A\times\mathscr{Q}^{\star}\}}$. Note that $(N_{u}([0,t]))_{t\geq 0}$ is a Poisson process with parameter $\Lambda^{a}(\mathscr{Q}^{\star})$.
We first deal with the initial particle $\varnothing$. The birth-time $\mathfrak{b}^{a}_{\varnothing}$ of particle $\varnothing$ equals 0. Enumerating the points of $\mathscr{P}_{\varnothing}$ by increasing first coordinate we obtain a sequence $(T_{\varnothing,j},\Delta_{\varnothing,j})_{j\geq 1}$ and note that $(\Delta_{\varnothing,j})_{j\geq 1}$ is a family of i.i.d $\mathscr{Q}^{\star}$-valued random variables with law $\Lambda^{a,\star}$. The $i$-th term of the sequence $\Delta_{\varnothing,j}$ is denoted by $\Delta_{\varnothing,j}(i)$. For any $t\geq 0$ we set,
$Y^{a}_{\varnothing}(t):=\xi_{\varnothing}(t)+\sum_{j=1}^{N_{\varnothing}^{0}([0,t])}\Delta_{\varnothing,j}(1),$ (47)
thus giving the generation 0. Now, let $n\geq 0$, assume that $\{(\mathfrak{b}^{a}_{u},Y^{a}_{u});\;|u|\leq n\}$ has been built, where $\mathfrak{b}^{a}_{u}$ denotes the birth-time of particle $u$. We now build generation $n+1$. For $u$ in generation $n$, enumerating the points of $\mathscr{P}_{u}$, restricted to $\mathfrak{b}^{a}_{u},\infty)\times\mathscr{Q}^{\star}$, by increasing first coordinate we obtain a sequence $(T_{u,j},\Delta_{u,j})_{j\geq 1}$. Again, $(\Delta_{u,j})_{j\geq 1}$ is a family of i.i.d $\mathscr{Q}^{\star}$-valued random variables with law $\Lambda^{a,\star}$. The $i$-th term of the sequence $\Delta_{u,j}$ is denoted by $\Delta_{u,j}(i)$.
Following [26], we divide $(\Delta_{u,j})_{j\geq 1}$ into (disjoint) classes corresponding to the truncation level of the children they represent. For that, let $q$ be a positive integer and define a new family $\Delta^{q}_{u,j}:=(\Delta^{q}_{u,j}(l))_{l\geq 1}$ as follows: $\Delta^{q}_{u,j}=\emptyset$ if the set $\{i\geq 2;\;\Delta_{u,j}(i)\in[q-1,q)\}$ is empty. Otherwise, the family $\Delta^{q}_{u,j}$ is made up of all the $\Delta_{u,j}(i)$, for $i\geq 2$, such that $\Delta_{u,j}(i)\in[q-1,q)$. By definition, $\Delta^{q}_{u,j}$ contains a finite number of elements that we rearrange in an increasing order. In order to see $\Delta^{q}_{u,j}$ as an element of $\mathscr{Q}$, we fill its tail with $\infty$. One can see that $\Delta^{q}_{u,j}=\emptyset$ for all $q>a$. By doing this, we will be able to index the branchings of all particles and that indexing will remain as $a$ gets larger, thus ensuring consistency when $a$ goes to infinity (see Definition A.4).
@@
Let $\bm{p}=(q,k,i)\in(\mathbb{N}^{*})^{3}$. The birth-time $\mathfrak{b}^{a}_{u\bm{p}}$ of the child $u\bm{p}$ of particle $u$ is set to be $\infty$ if $\mathfrak{b}^{a}_{u}=\infty$, $B^{q}_{u}(k)=\infty$ or $\Delta^{q}_{u,B^{q}_{u}(k)}(i)=\infty$ and we set $Y^{a}_{u\bm{p}}=\infty$. In particular, $Y^{a}_{u\bm{p}}=\infty$ when $q>a$. Otherwise, it is given by
$\mathfrak{b}^{a}_{u\bm{p}}:=T_{u,B^{q}_{u}(k)}.$
@@
$Y^{a}_{u\bm{p}}(t):=Y^{a}_{u}\big(\mathfrak{b}^{a}_{u\bm{p}}-\big)+\Delta^{q}_{u,B^{q}_{u}(k)}(i)+\xi_{u\bm{p}}(t-\mathfrak{b}^{a}_{u\bm{p}})+\sum_{j=1}^{N_{u\bm{p}}([\mathfrak{b}^{a}_{u\bm{p}},t])}\Delta_{u\bm{p},j}(1).$ (48)
@@
Let $u=u_{1}\cdots u_{\mathfrak{j}}\in\mathscr{U}$ for some $\mathfrak{j}\geq 1$. We extend the definition of $Y^{a}_{u}$ on $0,\mathfrak{b}^{a}_{u})$ in the following way: if $\mathfrak{j}\geq 2$, then for any $i\in\{1,\ldots,\mathfrak{j}-1\}$, $Y^{a}_{u}(t)=Y^{a}_{u_{1}\ldots u_{\mathfrak{i}}}(t)$ if $t\in[\mathfrak{b}^{a}_{u_{1}\cdots u_{\mathfrak{i}}},\mathfrak{b}^{a}_{u_{1}\cdots u_{\mathfrak{i}+1}})$ and $Y^{a}_{u}(t)=Y^{a}_{\varnothing}(t)$ if $t\in[0,\mathfrak{b}^{a}_{u_{1}})$. Otherwise, that is $\mathfrak{j}=1$, we set $Y^{a}_{u}(t)=Y^{q}_{\varnothing}(t)$ for any $t\in[0,\mathfrak{b}^{a}_{u})$. In other words, if the particle $u$ is not born yet at time $t$, then $Y^{a}_{u}(t)$ stands for the position at time $t$ of the most recent ancestor alive (at time $t$).
@@
$Y^{a}$ is referred to as a $a$-truncated branching subordinator. By a slight adaptation of Theorem 1.1(ii) in [14] we get that, for any $\lambda\geq 0$, the identity (13) holds true for $Y^{a}$ and $Y^{a}$ is a branching subordinator with finite birth intensity and characteristics $(d,\Lambda^{a})$ in the sense of Section 4.2 of [14].
@@
the canonical trajectory of particle $u$. Let $\bm{\xi}^{a}_{u}$, $u\in\mathscr{U}$ such that $\mathfrak{b}^{a}_{u}=\infty$, be i.i.d copies of $Y^{a}_{\varnothing}$ and independent of $(\bm{\xi}^{a}_{u};\;u\in\mathscr{U},\;\mathfrak{b}^{a}_{u}<\infty)$. Hence it is easily seen from (47) and (48) that, for all $a>a_{0}(\Lambda)$, $(\bm{\xi}^{a}_{u})_{u\in\mathscr{U}}$ is a collection of i.i.d subordinators with Laplace exponent $\phi$ and starting position 0.
@@
There exists a sequence $\{Y^{a};a\geq\lfloor a_{0}(\Lambda)\rfloor+1\}$ such that, for any $a\geq\lfloor a_{0}(\Lambda)\rfloor+1$, $Y^{a}$ is a branching subordinator with finite birth intensity $(d,\Lambda^{a})$ and, for any integer $a^{\prime}\geq a$
@@
We refer to Lemma 2.3 in [26] for a proof. For any $t\geq 0$, $\{Y^{a}(t);a\geq\lfloor a_{0}(\Lambda)\rfloor+1\}$ is an non-decreasing sequence of elements of $\mathscr{Q}$, which we recall is a subset of the space of point measures on $0,\infty)$.
@@
We now prove a technical lemma that will allow us to define without ambiguity the leftmost particle within some sets.
@@
Let $\Lambda$ be a measure on $\mathscr{Q}$ satisfying (3), (4) and Assumption 1 for some $\alpha\in(0,1)$. Let also $a>a_{0}(\Lambda)$. Then, for any fixed $t>0$, in the branching subordinator $Y^{a}$ with characteristic couple $(0,\Lambda^{a})$, we have almost surely that there are finitely many particles in the system at time $t$ and their positions are distinct. As a consequence, the set of times where the positions of particles are not distinct has null Lebesgue measure.